OpenStudy (anonymous):

how would i solve 8.34x=25.02 i dont understand how i would solve that equation using dividing

OpenStudy (anonymous):

Divide both sides by 8.34. This yields x = 3.

OpenStudy (anonymous):

Dividing 8.34x by 8.34 yields x. Dividing 25.02 by 8.34 yields 3. Do you understand now?

OpenStudy (anonymous):

yes thank you jhenahan and myininaya

myininaya (myininaya):

\[\frac{834}{100}x=\frac{2502}{100}\] \[\frac{100}{834}*\frac{834}{100}x=\frac{2502}{100}*\frac{100}{834}\] \[x=\frac{2502}{834}\] which can be simplified

OpenStudy (anonymous):

ok i get it now

myininaya (myininaya):

\[x=\frac{2502}{834}=\frac{834*3}{834}=3\]